# Measuring Chain of Thought Faithfulness by Unlearning Reasoning Steps

### v3

@@ 6.4 Quantifying Step Level Faithfulness
To quantitatively assess whether ff-soft identifies plausible steps as relevant, we conduct a user study on a random sample of $100$ instances. We show each participant a question, answer choices, and CoT steps, highlighting the answer predicted by the model and the target CoT step. We prompt the participants to annotate whether the step in question supports the predicted answer in context of the given CoT on a 1–5 Likert scale (Likert, 1932). We provide more details of the user study, data selection and the protocol in Appendix˜J.
We find a weak Pearson correlation of $0.15$ between ff-soft and human ratings of supportiveness. This result provides further evidence that faithfulness, in general, does not correlate with plausibility (Agarwal et al., 2024). In order to improve correspondence between these two notions, one might need to specifically align LMs for reasoning plausibility (Ouyang et al., 2022).
@@ 7 Discussion and Future Outlook
Recent works have turned to exploring test-time scaling via reasoning language models such as OpenAI o1 (Jaech et al., 2024) and DeepSeek R1 (DeepSeek-AI et al., 2025). Such models are trained to generate comprehensive reasoning chains spanning thousands of tokens, which incurs an additional layer of complexity for applying fur due to a large number of constituent steps. While a body of work strives to reduce overthinking in CoTs (Jin et al., 2024; Hassid et al., 2025; Xu et al., 2025; Amiri et al., 2025), the problem of length still persists in cases where long CoTs are necessary. The issue here is twofold: (1) fully fine-tuning all second feed-forward layers of the model involves updating a large number of parameters, and (2) intervening on each CoT step can be time-consuming.
@@
fur can innately be applied to individual CoT steps in parallel. However, not all CoT steps are equally salient for the prediction. We envision that models similar to verifiers (Lightman et al., 2024; Chen et al., 2024a; Jacovi et al., 2024) can detect, and rank, CoT steps most important for the models’ reasoning (thought anchors; Bogdan et al., 2025), which can then be prioritized for erasure. Recent work supporting our vision delves into analyzing the importance of, as well as controlling LMs chains-of-thought (Lee et al., 2025; Xiao et al., 2025; Yu et al., 2025). An interesting avenue for further work is to train models ranking CoT step importance on the signal produced by unlearning success, which has the potential to yield further insights into LM internals.
@@
To address the issue of space, we explore using LoRA (Hu et al., 2022) to reduce the number of active parameters. We apply LoRA tuning with ranks $8$, $32$ and $128$ to the FF2 matrix of two LLaMA models (3B and 8B) and $30$ instances from Sports and StrategyQA. As seen in Figure˜5, LoRA offers a potential alternative to full fine-tuning. Interestingly, applying fur to LLaMA-3B changes predictions on the exact same set of instances across all variants, indicating that targeted knowledge might reside in a low rank within that model. We detail the experimental setup and provide additional results in Appendix˜F.
@@ 8 Conclusion
Secondly, our approach relies on machine unlearning techniques, which are imperfect. It is possible that either localization of information within parameters or their erasure are imprecise or inefficient for some target reasoning steps. We rely on the rapid development of the field of machine unlearning and model editing to produce better and more precise methods such as CRISP Ashuach et al. (2025a) and PISCES (Gur-Arieh et al., 2025), which can seamlessly be integrated into our framework. While our method identifies faithful explanations with high precision, its recall cannot be guaranteed due to either unsuccessful unlearning, unfaithful explanation or the existence of alternative explanations. Furthermore, applying machine unlearning requires the capability to fine-tune the target model, which makes FUR not applicable to closed API-based models. Despite this limitation, we strongly believe that credible faithfulness of natural language explanations such as CoTs requires parameter access and interventions.
Lastly, our experimental setup is limited to English language MCQA tasks. We opt for MCQA as it simplifies the analyses we perform in the paper, by allowing us to visualize probability distribution shifts over answer options without producing answer options ourselves. Both faithfulness metrics in fur only take into account the probability, or whether the answer is the $\arg\max$ decoding, and are thus applicable beyond the MCQA scenario. Applying our method to other tasks such as long-form generation can be done by assessing whether the direct answer changes after unlearning a reasoning step. We opt for natural language tasks as factual information is conceptually easier to unlearn compared to, e.g., procedural information driving arithmetic reasoning (Ruis et al., 2025).
@@ Appendix A Background on Machine Unlearning
Motivated by the need to erase sensitive information from machine learning models (Cao and Yang, 2015; Harding et al., 2019; Ippolito et al., 2023), the field of machine unlearning emerged as an efficient alternative to filter-then-retrain-based approaches (Neel et al., 2021; Jang et al., 2023; Eldan and Russinovich, 2023; Liu et al., 2024, inter alia). Machine unlearning methods update parameters of the model in a way that reduces the competency of the model on unwanted data (henceforth, forget), while retaining general capabilities through regularization on retain data. Such methods decrease token probabilities on the forget data by gradient-based approaches (Jang et al., 2023; Eldan and Russinovich, 2023; Gandikota et al., 2024; Li et al., 2024; Zhang et al., 2024b) or directly updating parameters (Meng et al., 2022, 2023; Wu et al., 2023; Ashuach et al., 2025b).
@@
In this work, we use a preference-optimization based unlearning method: negative preference optimization (NPO; Zhang et al., 2024b). The core idea underpinning NPO is rooted in gradient ascent. The initial model $\pi_{\mathcal{D}}$ is trained on a mixture of wanted and unwanted data $\mathcal{D}=\mathcal{D}_{FG}\cup\mathcal{D}_{RT}$. Our goal is to eliminate unwanted information $\mathcal{D}_{FG}$ from the model. Therefore, applying the reverse language modeling objective on the forget data:
@@
Applying gradient ascent in this way runs into two practical issues. Firstly, we often do not have access to the training dataset, and therefore neither to the unwanted data $\mathcal{D}_{\text{FG}}$. Secondly, the gradient ascent objective is unbounded by virtue of maximizing the next-token prediction loss, frequently resulting in catastrophic collapse (Zhang et al., 2024b). Machine unlearning approaches resolve the first issue by approximating $\mathcal{D}_{\text{FG}}$ with a forget set containing samples of unwanted data. NPO resolves the second issue by constraining that the policy (predictive distribution) of the unlearned model $\pi_{\theta}$ should not diverge too far from a reference model $\pi_{\text{ref}}$. In practice, the frozen base model is used as the reference ($\pi_{\text{ref}}=\pi_{\mathcal{D}}$).
@@
where $\beta>0$ is the inverse temperature. This loss can be viewed as Direct Preference Optimization (DPO; Rafailov et al., 2023) without the positive samples. In practice, along with the loss term in Eq. (7), NPO also constrains the KL divergence between the unlearned and reference models on retain data in order to guarantee fluency, similar to other works (Li et al., 2024; Gandikota et al., 2024):
@@ Appendix B Alternatives to Measuring (Un)Faithfulness
A number of methods have been proposed with the goal of measuring faithfulness of model reasoning (Lanham et al., 2023; Atanasova et al., 2023; Bentham et al., 2024; Siegel et al., 2024; Chen et al., 2024b; Madsen et al., 2024, inter alia). However, not all of them are applicable to our setup. Some of the methods are designed for NLI tasks (Atanasova et al., 2023; Siegel et al., 2024; Parcalabescu and Frank, 2024) while others aim at multi-hop QA tasks (Parcalabescu and Frank, 2024; Lanham et al., 2023; Chen et al., 2024b; Bentham et al., 2024). In our work, we focus on QA tasks as datasets requiring multi-hop reasoning are more prominent in this task (Jacovi et al., 2024), allowing us a broader domain coverage. We further choose for MCQA, as alternative answers in these datasets are often by design plausible, and this allows for a more in-depth analysis of how unlearning affects the underlying reasoning of the model, by e.g. making it opt for plausible alternatives. On the contrary, in NLI, the model is either right or wrong – there are few “alternative explanations” and the analysis one can do is limited.
@@
We have considered other baselines applicable to CoT reasoning (cf. Table 1. in Zaman and Srivastava, 2025). Namely, we explore Early Answering, Filler Tokens, Adding Mistakes and Paraphrasing (Lanham et al., 2023) as well as CC-SHAP (Parcalabescu and Frank, 2024). Adding Mistakes is a simple contextual faithfulness method that works well, and we compare to its results in Table˜2. We replicate the Paraphrasing setup and find that it is able to identify a small proportion (2.842.84%, on average) of instances as unfaithful. See full results in Table˜4. Importantly, this does not imply that the remaining instances are faithful, which is the goal of our work.
@@
Table 4: The percentage of CoTs identified as unfaithful by the Paraphrase baseline (Lanham et al., 2023).
The remaining methods from Lanham et al. (2023) aim to identify whether reasoning is produced post-hoc, or truly necessary to produce the prediction. Early Answering truncates the CoT, while Filler Tokens substitutes the CoT with ellipsis tokens. Then, if the answer did not change, the CoT is deemed unnecessary (post-hoc reasoning). These measures do not aim to determine faithfulness of CoTs. Post-hoc reasoning can still be a true verbalization of latent reasoning, which is what both fur and Add-Mistake find Table˜2 since we only evaluate faithfulness cases where CoT and no-CoT predictions of models agree. Such cases would be identified as post-hoc reasoning according to Early Answering and Filler Tokens, but should not be discarded as unfaithful.
Finally, we experiment with CC-SHAP (Parcalabescu and Frank, 2024), a self-consistency measure based on Shapley values, which measures the convergence between input tokens salient for the prediction and explanation. We use the official implementation from the authors [^9] https://github.com/Heidelberg-NLP/CC-SHAP, but when applying SHAP to instances from our dataset, relative importances of tokens from input are frequently exactly zero (importances for reasoning do not behave in this manner), which results in NaN CC-SHAP scores in 90.2490.24% of instances across datasets and models. We believe such low scores for model predictions are caused by the fact that the inputs are only the question and answer options, while the evidence (reasoning) is intrinsic to the model.
@@ Appendix C Dataset and Model Statistics
To compute model predictions, we use letter completion. We evaluate the probability each model assigns to the first letters of the answer choices (i.e. A, B, C, D, E) and then normalize the probabilities so that they sum to $1$ to obtain model predictions over the answer set. We account for the verbosity issues raised by Wang et al. (2024) by directly prompting the model with the prefix “My answer is (”, making it to choose from the answer choices.
@@ Appendix D MCQA Task Prompts
We use two flavors of prompts when producing model predictions and the CoT for the evaluated tasks. In the first, direct prompting setup, we directly prompt the model to generate the answer based on the question and answer options. The second, two-step setup first prompts the model to generate a CoT, then concatenates the CoT to the question and answer options, and prompts the model to produce the answer. Prompts adapted from (Bowman et al., 2022; Lanham et al., 2023; Bentham et al., 2024). We conduct both prompting setups in zero-shot manner.
@@ Appendix E Unlearning Setup & Hyperparameters
We adapt the implementation of NPO+KL from the official repository. [^10] https://github.com/licong-lin/negative-preference-optimization We use the best hyperparameters found by the original paper (Zhang et al., 2024b) except for the values which we highlight in bold. See Table˜7 for values.
@@
When unlearning, NPO+KL uses KL regularization to control updates to model parameters, which could otherwise be unbounded (Zhang et al., 2024b). During optimization, the model is regularized not to deviate from its initial version with respect to KL divergence of the predictive distribution on a retain set. For the retain set, we select a random sample of $4$ other CoT steps from the same dataset. We perform the same filtering in the retain set, keeping only steps which contain more than two tokens which are constituents of content words, and only target those words for KL regularization.
@@ Appendix G Add-mistake Implementation
In this section, we detail our reimplementation of the add-mistake contextual faithfulness method (Lanham et al., 2023). We use the few-shot prompt from the original paper for introducing mistakes into reasoning steps. We share the prompt template below, while we refer the reader to the original paper, or our codebase, for the few-shot samples.
@@ Appendix H LLM-as-a-judge Setup
In order to evaluate whether the reasoning chains pre- and post-unlearning truly support different answer options, we follow the LLM-as-a-judge paradigm (Zheng et al., 2023a), leveraging gpt-4o-mini-2024-07-18 as the judge LM. We show the prompt we use below:
@@ Appendix I Additional Insights
We present each participant with annotation guidelines detailing the objective of the annotation, instructions detailing which aspects to pay attention to, and two annotation examples. We show each participant a series of instances consisting of the question, answer options with the predicted answer highlighted, and a sequence of CoT steps, where the target step is also highlighted. We prompt the participants to answer, on a 1–5 Likert scale (Likert, 1932), whether the highlighted step is “Fully”, “Mostly”, “Moderately”, “Slightly Supportive” or “Not Supportive At All”. We provide a screenshot from the annotation form in Figure˜11.

### v4

@@ 6.4 Quantifying Step Level Faithfulness
To quantitatively assess whether ff-soft identifies plausible steps as relevant, we conduct a user study on a random sample of $100$ instances. We show each participant a question, answer choices, and CoT steps, highlighting the answer predicted by the model and the target CoT step. We prompt the participants to annotate whether the step in question supports the predicted answer in context of the given CoT on a 1–5 Likert scale (likert1932technique). We provide more details of the user study, data selection and the protocol in Appendix˜J.
We find a weak Pearson correlation of $0.15$ between ff-soft and human ratings of supportiveness. This result provides further evidence that faithfulness, in general, does not correlate with plausibility (agarwal2024unrealiability). In order to improve correspondence between these two notions, one might need to specifically align LMs for reasoning plausibility (ouyang2022training).
@@ 7 Discussion and Future Outlook
Recent works have turned to exploring test-time scaling via reasoning language models such as OpenAI o1 (openai2024o1) and DeepSeek R1 (deepseek2025r1). Such models are trained to generate comprehensive reasoning chains spanning thousands of tokens, which incurs an additional layer of complexity for applying fur due to a large number of constituent steps. While a body of work strives to reduce overthinking in CoTs (jin2025impact; hassid2025overthink; wu2025chain; amiri2025lower), the problem of length still persists in cases where long CoTs are necessary. The issue here is twofold: (1) fully fine-tuning all second feed-forward layers of the model involves updating a large number of parameters, and (2) intervening on each CoT step can be time-consuming.
@@
fur can innately be applied to individual CoT steps in parallel. However, not all CoT steps are equally salient for the prediction. We envision that models similar to verifiers (lightman2024verify; chen2024process; jacovi2024chain) can detect, and rank, CoT steps most important for the models’ reasoning (thought anchors; bogdan2025anchors), which can then be prioritized for erasure. Recent work supporting our vision delves into analyzing the importance of, as well as controlling LMs chains-of-thought (lee2025encyclopedia; xiao2025limopro; yu2025explainable). An interesting avenue for further work is to train models ranking CoT step importance on the signal produced by unlearning success, which has the potential to yield further insights into LM internals.
@@
To address the issue of space, we explore using LoRA (hu2022lora) to reduce the number of active parameters. We apply LoRA tuning with ranks $8$, $32$ and $128$ to the FF2 matrix of two LLaMA models (3B and 8B) and $30$ instances from Sports and StrategyQA. As seen in Figure˜5, LoRA offers a potential alternative to full fine-tuning. Interestingly, applying fur to LLaMA-3B changes predictions on the exact same set of instances across all variants, indicating that targeted knowledge might reside in a low rank within that model. We detail the experimental setup and provide additional results in Appendix˜F.
@@ 8 Conclusion
Secondly, our approach relies on machine unlearning techniques, which are imperfect. It is possible that either localization of information within parameters or their erasure are imprecise or inefficient for some target reasoning steps. We rely on the rapid development of the field of machine unlearning and model editing to produce better and more precise methods such as CRISP ashuach2025crisp and PISCES (gur2025precise), which can seamlessly be integrated into our framework. While our method identifies faithful explanations with high precision, its recall cannot be guaranteed due to either unsuccessful unlearning, unfaithful explanation or the existence of alternative explanations. Furthermore, applying machine unlearning requires the capability to fine-tune the target model, which makes FUR not applicable to closed API-based models. Despite this limitation, we strongly believe that credible faithfulness of natural language explanations such as CoTs requires parameter access and interventions.
Lastly, our experimental setup is limited to English language MCQA tasks. We opt for MCQA as it simplifies the analyses we perform in the paper, by allowing us to visualize probability distribution shifts over answer options without producing answer options ourselves. Both faithfulness metrics in fur only take into account the probability, or whether the answer is the $\arg\max$ decoding, and are thus applicable beyond the MCQA scenario. Applying our method to other tasks such as long-form generation can be done by assessing whether the direct answer changes after unlearning a reasoning step. We opt for natural language tasks as factual information is conceptually easier to unlearn compared to, e.g., procedural information driving arithmetic reasoning (ruis2025procedural).
@@ Appendix A Background on Machine Unlearning
Motivated by the need to erase sensitive information from machine learning models (cao2015forget; harding2019understanding; ippolito2023preventing), the field of machine unlearning emerged as an efficient alternative to filter-then-retrain-based approaches (neel2021descent; jang2023unlearning; eldan2023harry; liu2024rethinking, inter alia). Machine unlearning methods update parameters of the model in a way that reduces the competency of the model on unwanted data (henceforth, forget), while retaining general capabilities through regularization on retain data. Such methods decrease token probabilities on the forget data by gradient-based approaches (jang2023unlearning; eldan2023harry; gandikota2024erasing; li2024wmdp; zhang2024npo) or directly updating parameters (meng2022locating; meng2023memit; wu2023depn; ashuach2024revs).
@@
In this work, we use a preference-optimization based unlearning method: negative preference optimization (NPO; zhang2024npo). The core idea underpinning NPO is rooted in gradient ascent. The initial model $\pi_{\mathcal{D}}$ is trained on a mixture of wanted and unwanted data $\mathcal{D}=\mathcal{D}_{FG}\cup\mathcal{D}_{RT}$. Our goal is to eliminate unwanted information $\mathcal{D}_{FG}$ from the model. Therefore, applying the reverse language modeling objective on the forget data:
@@
Applying gradient ascent in this way runs into two practical issues. Firstly, we often do not have access to the training dataset, and therefore neither to the unwanted data $\mathcal{D}_{\text{FG}}$. Secondly, the gradient ascent objective is unbounded by virtue of maximizing the next-token prediction loss, frequently resulting in catastrophic collapse (zhang2024npo). Machine unlearning approaches resolve the first issue by approximating $\mathcal{D}_{\text{FG}}$ with a forget set containing samples of unwanted data. NPO resolves the second issue by constraining that the policy (predictive distribution) of the unlearned model $\pi_{\theta}$ should not diverge too far from a reference model $\pi_{\text{ref}}$. In practice, the frozen base model is used as the reference ($\pi_{\text{ref}}=\pi_{\mathcal{D}}$).
@@
where $\beta>0$ is the inverse temperature. This loss can be viewed as Direct Preference Optimization (DPO; DBLP:conf/nips/RafailovSMMEF23) without the positive samples. In practice, along with the loss term in Eq. (7), NPO also constrains the KL divergence between the unlearned and reference models on retain data in order to guarantee fluency, similar to other works (li2024wmdp; gandikota2024erasing):
@@ Appendix B Alternatives to Measuring (Un)Faithfulness
A number of methods have been proposed with the goal of measuring faithfulness of model reasoning (lanham2023measuring; atanasova2023faithfulness; DBLP:journals/tmlr/BenthamSM24; siegel2024probabilities; chen2024counterfactual; madsen2024self, inter alia). However, not all of them are applicable to our setup. Some of the methods are designed for NLI tasks (atanasova2023faithfulness; siegel2024probabilities; parcalabescu2023measuring) while others aim at multi-hop QA tasks (parcalabescu2023measuring; lanham2023measuring; chen2024counterfactual; DBLP:journals/tmlr/BenthamSM24). In our work, we focus on QA tasks as datasets requiring multi-hop reasoning are more prominent in this task (jacovi2024chain), allowing us a broader domain coverage. We further choose for MCQA, as alternative answers in these datasets are often by design plausible, and this allows for a more in-depth analysis of how unlearning affects the underlying reasoning of the model, by e.g. making it opt for plausible alternatives. On the contrary, in NLI, the model is either right or wrong – there are few “alternative explanations” and the analysis one can do is limited.
@@
We have considered other baselines applicable to CoT reasoning (cf. Table 1. in zaman2025causal). Namely, we explore Early Answering, Filler Tokens, Adding Mistakes and Paraphrasing (lanham2023measuring) as well as CC-SHAP (parcalabescu2023measuring). Adding Mistakes is a simple contextual faithfulness method that works well, and we compare to its results in Table˜2. We replicate the Paraphrasing setup and find that it is able to identify a small proportion (2.842.84%, on average) of instances as unfaithful. See full results in Table˜4. Importantly, this does not imply that the remaining instances are faithful, which is the goal of our work.
@@
Table 4: The percentage of CoTs identified as unfaithful by the Paraphrase baseline (lanham2023measuring).
The remaining methods from lanham2023measuring aim to identify whether reasoning is produced post-hoc, or truly necessary to produce the prediction. Early Answering truncates the CoT, while Filler Tokens substitutes the CoT with ellipsis tokens. Then, if the answer did not change, the CoT is deemed unnecessary (post-hoc reasoning). These measures do not aim to determine faithfulness of CoTs. Post-hoc reasoning can still be a true verbalization of latent reasoning, which is what both fur and Add-Mistake find Table˜2 since we only evaluate faithfulness cases where CoT and no-CoT predictions of models agree. Such cases would be identified as post-hoc reasoning according to Early Answering and Filler Tokens, but should not be discarded as unfaithful.
Finally, we experiment with CC-SHAP (parcalabescu2023measuring), a self-consistency measure based on Shapley values, which measures the convergence between input tokens salient for the prediction and explanation. We use the official implementation from the authors [^9] https://github.com/Heidelberg-NLP/CC-SHAP, but when applying SHAP to instances from our dataset, relative importances of tokens from input are frequently exactly zero (importances for reasoning do not behave in this manner), which results in NaN CC-SHAP scores in 90.2490.24% of instances across datasets and models. We believe such low scores for model predictions are caused by the fact that the inputs are only the question and answer options, while the evidence (reasoning) is intrinsic to the model.
@@ Appendix C Dataset and Model Statistics
To compute model predictions, we use letter completion. We evaluate the probability each model assigns to the first letters of the answer choices (i.e. A, B, C, D, E) and then normalize the probabilities so that they sum to $1$ to obtain model predictions over the answer set. We account for the verbosity issues raised by wang2024letter by directly prompting the model with the prefix “My answer is (”, making it to choose from the answer choices.
@@ Appendix D MCQA Task Prompts
We use two flavors of prompts when producing model predictions and the CoT for the evaluated tasks. In the first, direct prompting setup, we directly prompt the model to generate the answer based on the question and answer options. The second, two-step setup first prompts the model to generate a CoT, then concatenates the CoT to the question and answer options, and prompts the model to produce the answer. Prompts adapted from (bowman2022scalable; lanham2023measuring; DBLP:journals/tmlr/BenthamSM24). We conduct both prompting setups in zero-shot manner.
@@ Appendix E Unlearning Setup & Hyperparameters
We adapt the implementation of NPO+KL from the official repository. [^10] https://github.com/licong-lin/negative-preference-optimization We use the best hyperparameters found by the original paper (zhang2024npo) except for the values which we highlight in bold. See Table˜7 for values.
@@
When unlearning, NPO+KL uses KL regularization to control updates to model parameters, which could otherwise be unbounded (zhang2024npo). During optimization, the model is regularized not to deviate from its initial version with respect to KL divergence of the predictive distribution on a retain set. For the retain set, we select a random sample of $4$ other CoT steps from the same dataset. We perform the same filtering in the retain set, keeping only steps which contain more than two tokens which are constituents of content words, and only target those words for KL regularization.
@@ Appendix G Add-mistake Implementation
In this section, we detail our reimplementation of the add-mistake contextual faithfulness method (lanham2023measuring). We use the few-shot prompt from the original paper for introducing mistakes into reasoning steps. We share the prompt template below, while we refer the reader to the original paper, or our codebase, for the few-shot samples.
@@ Appendix H LLM-as-a-judge Setup
In order to evaluate whether the reasoning chains pre- and post-unlearning truly support different answer options, we follow the LLM-as-a-judge paradigm (zheng2023llmasjudge), leveraging gpt-4o-mini-2024-07-18 as the judge LM. We show the prompt we use below:
@@ Appendix I Additional Insights
We present each participant with annotation guidelines detailing the objective of the annotation, instructions detailing which aspects to pay attention to, and two annotation examples. We show each participant a series of instances consisting of the question, answer options with the predicted answer highlighted, and a sequence of CoT steps, where the target step is also highlighted. We prompt the participants to answer, on a 1–5 Likert scale (likert1932technique), whether the highlighted step is “Fully”, “Mostly”, “Moderately”, “Slightly Supportive” or “Not Supportive At All”. We provide a screenshot from the annotation form in Figure˜11.